A wire , 20m long is heated from a temperature of -5⁰c to 55⁰c,if the change in length is 0.020m, calculate the linear expansivity of the wire


1
Expert's answer
2022-01-23T16:13:10-0500

Given:

"l_0=20\:\rm m"

"\Delta l=0.020\:\rm m"

"\Delta t=60\:^\circ\rm C"


The change in length of a wire at termal expansion is given by

"\Delta l=l_0\alpha\Delta t"

Hence, the linear expansivity of the wire

"\alpha=\frac{\Delta l}{l_0\Delta t}=\frac{0.020}{20*60}=1.7*10^{-5}\:\rm K^{-1}"